OpenStudy (anonymous):

After a discount of 20%, the price of fitness equipment, including tax, was $73.53. The following equation can be used to find x, the price of the equipment without the discount: x - 0.2x + 0.07(x - 0.2x) = 73.53 What is the price of the fitness equipment without the discount? $74.39 $79.06 $85.90 $95.49

OpenStudy (anonymous):

(x - 0.2x) (1 + 0.07) = 73.53 0.8x (1.07) = 73.53 0.856x = 73.53 x = 73.53 / 0.856 = 85.90

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

medal?